# PoseAnimate: Zero-shot high fidelity pose controllable character animation

### v1

@@ Abstract
Image-to-video(I2V) generation aims to create a video sequence from a single image, which requires high temporal coherence and visual fidelity with the source image. However, existing approaches suffer from character appearance inconsistency and poor preservation of fine details. Moreover, they require a large amount of video data for training, which can be computationally demanding. To address these limitations, we propose PoseAnimate, a novel zero-shot I2V framework for character animation. PoseAnimate contains three key components: 1) Pose-Aware Control Module (PACM) incorporates diverse pose signals into conditional embeddings, to preserve character-independent content and maintain precise alignment of actions. 2) Dual Consistency Attention Module (DCAM) enhances temporal consistency, and retains character identity and intricate background details. 3) Mask-Guided Decoupling Module (MGDM) refines distinct feature perception, improving animation fidelity by decoupling the character and background. We also propose a Pose Alignment Transition Algorithm (PATA) to ensure smooth action transition. Extensive experiment results demonstrate that our approach outperforms the state-of-the-art training-based methods in terms of character consistency and detail fidelity. Moreover, it maintains a high level of temporal coherence throughout the generated animations.
{strip}[Uncaptioned image]
Figure 1: Our PoseAnimate framework is capable of generating smooth and high-quality character animations for character images across various pose sequences.
@@ 1 Introduction
Image animation siarohin2019first ; siarohin2019animating ; siarohin2021motion ; wang2022latent ; zhao2022thin is a task that brings life into static images by seamlessly transforming them into dynamic and realistic videos. It involves the transformation of still images into a sequence of frames that exhibit smooth and coherent motions. In this task, character animation has gained significant attention due to its valuable applications in various scenarios, such as television production, game development, online retail and artistic creation, etc. However, minor motion variations hardly meet with the requirements. The goal of character animation is to make the character in image perform target pose sequences, while maintaining identity consistency and visual coherence. In early works, most of character animation were driven by traditional animation techniques, which involves meticulous frame-by-frame drawing or manipulation. In the subsequent era of deep learning, the advent of generative models goodfellow2014generative ; zhu2017unpaired ; karras2019style drove the shift towards data-driven and automated approaches ren2020deep ; chan2019everybody ; zhang2022exploring . However, there are still ongoing challenges in achieving highly realistic and visually consistent animations, especially when dealing with complex motions, fine-grained details, and long-term temporal coherence.
@@
Recently, diffusion models ho2020denoising have demonstrated groundbreaking generative capabilities. Driven by the open source text-to-image diffusion model Stable Diffusion rombach2022high , the realm of video generation has achieved unprecedented progress in terms of visual quality and content richness. Hence, several endeavors wang2023disco ; xu2023magicanimate ; hu2023animate have sought to extrapolate the text-to-video(T2V) methods to image-to-video(I2V) by training additional image feature preserving networks and adapt them to the task of character animation. Nevertheless, these training-based methods do not possess accurate feature preservation capabilities for arbitrary open-domain images, and suffer from notable deficiencies in appearance control and loss of fine details. Furthermore, they require additional training data and computational overhead.
@@
To this end, we contemplate employing a more refined and efficient resolution, image reconstruction for feature preservation, to tackle this problem. We propose PoseAnimate, depicted in Fig. 2, a zero-shot reconstruction-based I2V framework for pose controllable character animation video generation. PoseAnimate introduces a pose-aware control module(PACM), shown in Fig. 3 which optimizes the text embedding twice based on the original and target pose conditions respectively finally resulting a unique pose-aware embedding for each generated frame. This optimization strategy allows for the generated actions aligned to the target pose while contributing to keep the character-independent scene consistent. However, the introduction of a new target pose in the second optimization, which differs from the original pose, inevitably undermines the reconstruction of the character’s identity and background. Thus, we further devise a dual consistency attention module(DCAM), as dedicated in the right part of Fig. 2, to address the disruption, in addition to maintain a smooth temporal progression. Since directly employing the entire attention map or key for attention fusion may result in the loss of fine-grained detail perception. We propose a mask-guided decoupling module(MGDM) to enable independent and focused spatial attention fusion for both the character and background. As such, our framework promises to capture the intricate character and background details, thereby effectively enhancing the fidelity of the animation. Besides, for the sake of adaptation to various scales and positions of target pose sequences, a pose alignment transition algorithm(PATA) is designed to ensure pose alignment and smooth transitions. Through combination of these novel modules, PoseAnimate achieves promising character animation results, as shown in Fig. 1, in a more efficient manner with lower computational overhead.
@@
To summarize, our contributions are as follows: 1) We pioneer a reconstruction-based approach to handle the task of character animation and propose PoseAnimate, a novel zero-shot framework, which generates coherent high-quality videos for arbitrary character images under various pose sequences, without any training of the network. To the best of our knowledge, we are the first to explore a training-free approach to character animation. 2) We propose a pose-aware control module that enables precise alignment of actions while maintaining consistency across character-independent scenes. 3) We decouple the character and the background regions, performing independent inter-frame attention fusion for them, which significantly enhances visual fidelity. 4) Experiment results demonstrate the superiority of PoseAnimate compared with the state-of-the-art training-based methods in terms of character consistency and image fidelity.
@@ 2.1 Diffusion Models for Video Generation
Image generation has made significant progress due to the advancement of Diffusion Models(DMs) ho2020denoising . Motivated by DM-based image generation rombach2022high , some works yang2023diffusion ; ho2022imagen ; nikankin2022sinfusion ; esser2023structure ; blattmann2023align explore DMs for video generation. Most video generation methods incorporate temporal modules to pretrained image diffusion models, extending 2D U-Net to 3D U-Net. Recent works control the generation of videos with multiple conditions. For text-guided video generation, these works he2022latent ; ge2023preserve ; gu2023reuse usually tokenize text prompts with a pretrained image-language model, such as CLIP radford2021learning , to control video generation through cross-attention. Due to the imperfect alignment between language and visual modalities in existing image-language models, text-guided video generation can’t achieve high textual alignment. Alternative methods wang2023videocomposer ; chen2023videocrafter1 ; blattmann2023stable employ images as guidance for video generation. These works encode reference images to text token space, which benefits capturing visual semantic information. VideoComposerwang2023videocomposer combines textual conditions, spatial conditions(e.g., depth, sketch, reference image) and temporal conditions(e.g., motion vector) through Spatio-Temporal Condition encoders. VideoCrafter1chen2023videocrafter1 introduces a text-aligned rich image embedding to capture details both from text prompts and reference images. Stable Video Diffusion blattmann2023stable is a latent diffusion model for high-resolution T2V and I2V generation, which sets three different stages for training: text-to-image pretraining, video pretraining, and high-quality video finetuning.
@@ 2.2 Video Generation with Human Pose
Generating videos with human pose is currently a popular task. Compared to other conditions, human pose can better guide the synthesis of motions in videos, which ensures good temporal consistency. Follow your posema2023follow introduces a two-stage method to generate pose-controllable character videos. Many studies wang2023disco ; karras2023dreampose ; xu2023magicanimate ; hu2023animate try to generate character videos from still images via pose sequence, which needs to preserve consistency of appearance from source images as well. Inspired by ControlNetzhang2023adding , DisCowang2023disco realizes disentangled control of human foreground, background and pose, which enables faithful human video generation. To increase fidelity to the reference human images, DreamPosekarras2023dreampose proposes an adapter to models CLIP and VAE image embeddings. MagicAnimatexu2023magicanimate adopts ControlNetzhang2023adding to extract motion conditions. It also introduces a appearance encoder to model reference images embedding. Animate Anyonehu2023animate designs a ReferenceNet to extract detail features from reference images, combined with a pose guider to guarantee motion generation.
@@ 3 Method
Given a source character image $I_{s}$, and a desired pose sequence $P=\{p_{i}\}_{i=1}^{M}$, where $M$ is the length of sequence. In the generated animation, we adopt a progressive approach to transition the character seamlessly from the original pose $p_{s}$ to the desired pose sequence $P=\{p_{i}\}_{i=1}^{M}$. We first facilitate the Pose Alignment Transition Algorithm(PATA), derailed in supplementary material, to smoothly interpolate $t$ intermediate frames between the source pose $p_{s}$ and the target pose sequence $P=\{p_{i}\}_{i=1}^{M}$. Simultaneously, it aligns each target pose $p_{i}$ with the source pose $p_{s}$ to compensate for their discrepancies in terms of position and scale. As a result, the final target pose sequence is $P=\{p_{i}\}_{i=0}^{N}$, where $N=M+t$. It is worth noting that the first frame $x_{0}$ in our generated animation $X=\{x_{i}\}_{i=0}^{N}$ is identical to the source image $I_{s}$. Secondly, we propose a pose-aware control module(PACM) that optimizes a unique pose-aware embedding for each generated frame. This module can eliminate perturbation of original character posture, thereby ensuring the generated actions aligned with the target pose. Furthermore, it also maintains consistency of content irrelevant to characters. Thirdly, a dual consistency attention module(DCAM) is developed to ensure consistency of the character identity and improve temporal consistency. In addition, we design a mask-guided decoupling module(MGDM) to further enhance perception of characters and backgrounds details. The overview of our PoseAnimate is shown in Fig. 2.
In this section, we first give an introduction of Stable Diffusion in Sec 3.1. Subsequently, Sec 3.2 introduces the incorporation of motion awareness into pose-aware embedding. The proposed dual consistency control is elaborated in Sec 3.3, followed by mask-guided decoupling module in Sec 3.4.
@@ 3.1 Preliminaries on Stable Diffusion
Stable Diffusion rombach2022high has demonstrated strong text-to-image generation ability through a diffusion model in a latent space constructed by a pair of image encoder $\mathcal{E}$ and decoder $\mathcal{D}$. For an input image $\mathcal{I}$, the encoder $\mathcal{E}$ first maps it to a lower dimensional latent code $z_{0}=\mathcal{E}(\mathcal{I})$, then Gaussian noise is gradually added to $z_{0}$ through the diffusion forward process:
@@
where $t=1,...,T$, denotes the timesteps, $\beta_{t}\in(0,1)$ is a predefined noise schedule. Through a parameterization trick, we can directly sample $z_{t}$ from $z_{0}$:
@@
where $\bar{\alpha_{t}}=\prod_{i=1}^{t}\alpha_{i}$, and $\alpha_{t}=1-\beta_{t}$. Diffusion model uses a neural network $\epsilon_{\theta}$ to learn to predict the added noise $\epsilon$ by minimizing the mean square error of the predicted noise:
@@
where $\mathbf{c}$ is embedding of textual prompt. And we can adopt a deterministic sampling process song2020denoising to iteratively recover $z_{0}\sim\mathcal{P}_{data}(z)$ from random noise $z_{T}$:
@@ 3.2 Pose-Aware Control Module
For generating a high fidelity character animation from a source image, two tasks need to be accomplished. Firstly, it is critical to preserve the consistency of original character and background in generated animation. In contrast to other approaches karras2023dreampose ; xu2023magicanimate ; hu2023animate that rely on training additional spatial preservation networks for consistency identity, we achieve it through a computationally efficient reconstruction-based method. Secondly, the actions in generated frames needs to align with the target poses. Although the pre-trained OpenPose ControlNet zhang2023adding has great spatial control capabilities in controllable condition synthesis, our purpose is to discard the original pose and generate new continuous motion. Therefore, directly introducing pose signals through ControlNet may result in conflicts with the original pose, resulting in severe ghosting and blurring in motion areas.
@@
In light of this, we propose the pose-aware control module, as illustrated in the Fig. 3. Inspired by the idea of inversion in image editing mokady2023null , we achieve the perception of pose signals by optimizing the text embedding $\varnothing_{text}$ twice based on the original pose $p_{s}$ and target pose $p_{i}$ respectively. In the first optimization, i.e. pose-aware inversion, we iteratively refine the original text embedding $\varnothing_{text}$ to accurately reconstruct the intricate details of the source image $I_{s}$ under the original pose $p_{s}$. Building upon the optimized source embeddings $\{\varnothing_{s,t}\}_{t=1}^{T}$ obtained from this process, we then proceed with the second optimization, i.e. pose-aware embedding optimization, where we inject the target pose signals $P=\{p_{i}\}_{i=1}^{N}$ into the optimized pose-aware embeddings {{~$\{\{\widetilde{\varnothing}_{x_{i},t}\}_{t=1}^{T}\}_{i=1}^{N}$, as detailed in Alg. 1. Perceiving the target pose signals, these optimized pose-aware embeddings {{~$\{\{\widetilde{\varnothing}_{x_{i},t}\}_{t=1}^{T}\}_{i=1}^{N}$ ensure a flawless alignment between the generated character actions and the target poses, while upholding the consistency of character-independent content.
@@
Specifically, to incorporate the pose signals, we integrate ControlNet into all processes of the module. Diverging from null-text inversion mokady2023null that achieves image reconstruction by optimizing unconditional embeddings ho2022classifier , our pose-aware inversion optimizes the conditional embedding $\varnothing_{text}$ of text prompt $C$ during the reconstruction process. The motivation stems from the observation that conditional embedding contains more abundant and robust semantic information, which endows it with a heightened potential for encoding pose signals.
Figure 3: Illustration of pose-aware control module. We optimize the text embedding twice to inject motion awareness into pose-aware embedding.
@@
1:  Set guidance scale = 1.0. Calculate DDIM inversion latent code $Z_{0},...,Z_{T}$ corresponding to input image $I_{s}$.
2:  Set guidance scale = 7.5. Obtain optimized source embeddings $\{\varnothing_{s,t}\}_{t=1}^{T}$ through pose-aware inversion (Fig. 3).

### v3

# Zero-shot High-fidelity and Pose-controllable Character Animation

@@ Abstract
Image-to-video (I2V) generation aims to create a video sequence from a single image, which requires high temporal coherence and visual fidelity. However, existing approaches suffer from inconsistency of character appearances and poor preservation of fine details. Moreover, they require a large amount of video data for training, which can be computationally demanding. To address these limitations, we propose PoseAnimate, a novel zero-shot I2V framework for character animation. PoseAnimate contains three key components: 1) a Pose-Aware Control Module (PACM) that incorporates diverse pose signals into text embeddings, to preserve character-independent content and maintain precise alignment of actions. 2) a Dual Consistency Attention Module (DCAM) that enhances temporal consistency and retains character identity and intricate background details. 3) a Mask-Guided Decoupling Module (MGDM) that refines distinct feature perception abilities, improving animation fidelity by decoupling the character and background. We also propose a Pose Alignment Transition Algorithm (PATA) to ensure smooth action transition. Extensive experiment results demonstrate that our approach outperforms the state-of-the-art training-based methods in terms of character consistency and detail fidelity. Moreover, it maintains a high level of temporal coherence throughout the generated animations.
Figure 1: PoseAnimate framework is capable of generating smooth and high-quality character animations for static character images across various pose sequences.
@@ 1 Introduction
Image animation Siarohin et al. (2019b, a, 2021); Wang et al. (2022); Zhao and Zhang (2022) is a task that brings static images to life by seamlessly transforming them into dynamic and realistic videos. It involves the transformation of still images into a sequence of frames that exhibit smooth and coherent motions. In this task, character animation has gained significant attention due to its valuable applications in various scenarios, such as television production, game development, online retail and artistic creation, etc. However, minor motion variations hardly meet with the requirements. The goal of character animation is to make the character in the image perform target pose sequences, while maintaining identity consistency and visual coherence. In early works, most of character animation was driven by traditional animation techniques, which involves meticulous frame-by-frame drawing or manipulation. In the subsequent era of deep learning, the advent of generative models Goodfellow et al. (2014); Zhu et al. (2017); Karras et al. (2019) drove the shift towards data-driven and automated approaches Ren et al. (2020); Chan et al. (2019); Zhang et al. (2022). However, there are still ongoing challenges in achieving highly realistic and visually consistent animations, especially when dealing with complex motions, fine-grained details, and long-term temporal coherence.
@@
Recently, diffusion models Ho et al. (2020) have demonstrated groundbreaking generative capabilities. Driven by the open source text-to-image diffusion model Stable Diffusion Rombach et al. (2022), the realm of video generation Xing et al. (2023c) has achieved unprecedented progress in terms of visual quality and content richness. Hence, several endeavors Wang et al. (2023a); Xu et al. (2023); Hu et al. (2023) have sought to extrapolate the text-to-video (T2V) methods to image-to-video (I2V) by training additional image feature preserving networks and adapt them to the character animation task. Nevertheless, these training-based methods face challenges in accurately preserving features for arbitrary images and exhibit notable deficiencies in appearance control and loss of details. Additionally, they typically rely on extensive training data and significant computational overhead.
@@
To this end, we contemplate employing a more refined and efficient approach, image reconstruction for feature preservation, to tackle this problem. We propose PoseAnimate, depicted in Fig. 2, a zero-shot reconstruction-based I2V framework for pose controllable character animation. PoseAnimate introduces a pose-aware control module (PACM), shown in Fig. 3 which optimizes the text embedding twice based on the original and target pose conditions respectively, finally resulting a unique pose-aware embedding for each generated frame. This optimization strategy allows for the generated actions to be aligned with the target pose while keeping the character-independent scene consistent. However, the introduction of a new target pose in the second optimization, which differs from the original pose, inevitably undermines the reconstruction of the character identity and background. Thus, we further devise a dual consistency attention module (DCAM), as dedicated in the right part of Fig. 2, to address the disruption, in addition to maintain a smooth temporal progression. Since directly employing the entire attention map or key for attention fusion may result in loss of fine-grained detail perception. We propose a mask-guided decoupling module (MGDM) to enable independent and focused spatial attention fusion for both the character and background. As such, our framework is able to capture the intricate character and background details, thereby effectively enhancing the fidelity of the animation. In addition, for the sake of adaptation to various scales and positions of target pose sequences, a pose alignment transition algorithm (PATA) is designed to ensure pose alignment and smooth transitions. Through combination of these novel modules, PoseAnimate achieves promising character animation results, as shown in Fig. 1, in a more efficient manner with lower computational overhead.
@@
In summary, our contributions are as follows: (1) We introduce a reconstruction-based approach to handle the task of character animation and propose PoseAnimate, a novel zero-shot framework, which generates coherent high-quality videos for arbitrary character images under various pose sequences, without any training of the network. To the best of our knowledge, we are the first to explore a training-free approach to character animation. (2) We propose a pose-aware control module that enables precise alignment of actions while maintaining consistency across character-independent scenes. (3) We decouple the character and the background regions, performing independent inter-frame attention fusion for them, which significantly enhances visual fidelity. (4) Experiment results demonstrate the superiority of PoseAnimate compared with the state-of-the-art training-based methods in terms of character consistency and image fidelity.
@@ 2.1 Diffusion Models for Video Generation
Image generation has made significant progress due to the advancement of Diffusion Models (DMs) Ho et al. (2020). Motivated by DM-based image generation Rombach et al. (2022), some works Yang et al. (2023); Ho et al. (2022); Nikankin et al. (2022); Esser et al. (2023); Xing et al. (2023a); Blattmann et al. (2023b); Xing et al. (2023b) explore DMs for video generation. Most video generation methods incorporate temporal modules to pre-trained image diffusion models, extending 2D U-Net to 3D U-Net. Recent works control the generation of videos with multiple conditions. For text-guided video generation, these works He et al. (2022); Ge et al. (2023); Gu et al. (2023) usually tokenize text prompts with a pre-trained image-language model, such as CLIP Radford et al. (2021), to control video generation through cross-attention. Due to the imperfect alignment between language and visual modalities in existing image-language models, text-guided video generation cannot achieve high textual alignment. Alternative methods Wang et al. (2023b); Chen et al. (2023); Blattmann et al. (2023a) employ images as additional guidance for video generation. These works encode reference images to token space, helping capturing visual semantic information. VideoComposer Wang et al. (2023b) combines textual conditions, spatial conditions (e.g., depth, sketch, reference image) and temporal conditions (e.g., motion vector) through Spatio-Temporal Condition encoders. VideoCrafter1 Chen et al. (2023) introduces a text-aligned rich image embedding to capture details both from text prompts and reference images. Stable Video Diffusion Blattmann et al. (2023a) is a latent diffusion model for high-resolution T2V and I2V generation, which sets three different stages for training: text-to-image pretraining, video pretraining, and high-quality video finetuning.
@@ 2.2 Video Generation with Human Pose
Generating videos with human pose is currently a popular task. Compared to other conditions, human pose can better guide the synthesis of motions in videos, which ensures good temporal consistency. Follow your pose Ma et al. (2023) introduces a two-stage method to generate pose-controllable character videos. Many studies Wang et al. (2023a); Karras et al. (2023); Xu et al. (2023); Hu et al. (2023) try to generate character videos from still images via pose sequence, which needs to preserve consistency of appearance from source images as well. Inspired by ControlNet Zhang et al. (2023), DisCo Wang et al. (2023a) realizes disentangled control of human foreground, background and pose, which enables faithful human video generation. To increase fidelity to reference human images, DreamPose Karras et al. (2023) proposes an adapter to models CLIP and VAE image embeddings. MagicAnimate Xu et al. (2023) adopts ControlNet Zhang et al. (2023) to extract motion conditions. It also introduces a appearance encoder to model reference images embedding. Animate Anyone Hu et al. (2023) designs a ReferenceNet to extract detail features from reference images, combined with a pose guider to guarantee motion generation.
@@ 3 Method
Given a source character image $I_{s}$, and a desired pose sequence $P=\{p_{i}\}_{i=1}^{M}$, where $M$ is the length of the sequence. In the generated animation, we adopt a progressive approach to seamlessly transition the character from the source pose $p_{s}$ to the desired pose sequence $P=\{p_{i}\}_{i=1}^{M}$. We first facilitate the Pose Alignment Transition Algorithm (PATA) to smoothly interpolate $t$ intermediate frames between the source pose $p_{s}$ and the desired pose $P=\{p_{i}\}_{i=1}^{M}$. Simultaneously, it aligns each pose $p_{i}$ with the source pose $p_{s}$ to compensate for their discrepancies in terms of position and scale. As a result, the final target pose sequence is $P=\{p_{i}\}_{i=0}^{N}$, where $N=M+t$. It is worth noting that the first frame $x_{0}$ in our generated animation $X=\{x_{i}\}_{i=0}^{N}$ is identical to the source image $I_{s}$. Secondly, we propose a pose-aware control module (PACM) that optimizes a unique pose-aware embedding for each generated frame. This module can eliminate perturbation of the original character posture, thereby ensuring the generated actions aligned with the target pose $P$. Furthermore, it also maintains consistency of content irrelevant to characters. Thirdly, a dual consistency attention module (DCAM) is developed to ensure consistency of the character identity and improve temporal consistency. In addition, we design a mask-guided decoupling module (MGDM) to further enhance perception of character and background details. The overview of our PoseAnimate is depicted in Fig. 2.
In this section, we begin with a brief introduction to Stable Diffusion in Sec. 3.1. Subsequently, Sec. 3.2 introduces the incorporation of motion awareness into pose-aware embedding. The proposed dual consistency control module is elaborated in Sec. 3.3, followed by the mask-guided decoupling module in Sec. 3.4.
@@ 3.1 Preliminaries on Stable Diffusion
Stable Diffusion Rombach et al. (2022) has demonstrated strong text-to-image generation ability through a diffusion model in a latent space constructed by a pair of image encoder $\mathcal{E}$ and decoder $\mathcal{D}$. For an input image $\mathcal{I}$, the encoder $\mathcal{E}$ first maps it to a lower dimensional latent code $z_{0}=\mathcal{E}(\mathcal{I})$, then Gaussian noise is gradually added to $z_{0}$ through the diffusion forward process:
@@
where $t=1,...,T$, denotes the timesteps, $\beta_{t}\in(0,1)$ is a predefined noise schedule, and $\mathbf{I}$ is identity matrix. Through a parameterization trick, we can directly sample $z_{t}$ from $z_{0}$:
@@
where $\bar{\alpha_{t}}=\prod_{i=1}^{t}\alpha_{i}$, and $\alpha_{t}=1-\beta_{t}$. Diffusion models use a neural network $\epsilon_{\theta}$ to learn to predict the added noise $\epsilon$ by minimizing the mean square error of the predicted noise:
@@
where $\mathbf{c}$ is embedding of textual prompt. During inference, we can adopt a deterministic DDIM sampling Song et al. (2020), to iteratively recover a denoised representation $x_{0}$ from standard Gaussian noise $z_{T},z_{T}\sim\mathcal{N}(0,I)$:
@@
Then $x_{0}$ is decoded into an output image $\mathcal{I}^{{}^{\prime}}=\mathcal{D}(x_{0})$ using the pre-trained decoder $\mathcal{D}$.
@@ 3.2 Pose-Aware Control Module
For generating a high fidelity character animation from a static image, two tasks need to be accomplished. Firstly, it is critical to preserve the consistency of original character and background in generated animation. In contrast to other approaches Karras et al. (2023); Xu et al. (2023); Hu et al. (2023) that rely on training additional spatial preservation networks for consistency identity, we achieve it through a computationally efficient reconstruction-based method. Secondly, the actions in the generated frames need to align with the target poses. Although the pre-trained OpenPose ControlNet Zhang et al. (2023) has great spatial control capabilities in controllable condition synthesis, our purpose is to discard the original pose and generate new continuous motion. Therefore, directly introducing pose signals through ControlNet may result in conflicts with the original pose, resulting in severe ghosting and blurring in motion areas.
@@
In light of this, we propose the pose-aware control module, as illustrated in Fig. 3. Inspired by the idea of inversion in image editing Mokady et al. (2023), we achieve the perception of pose signals by optimizing the text embedding $\varnothing_{text}$ twice based on the source pose $p_{s}$ and the target pose $p_{i}$, respectively. In the first optimization, i.e. pose-aware inversion, we involve a progressive optimization process for $\varnothing_{text}$ to accurately reconstruct the source image $I_{s}$ under the source pose $p_{s}$. We initialize $\bar{Z}_{s,T}=Z_{T}$, $\varnothing_{s,T}=\varnothing_{text}$, and perform the following optimization for the timesteps $t=T,\ldots,1$, each step for $n$ inner iterations:
$\small\min_{\varnothing_{s,t}}\|{Z_{t-1}-z_{t-1}(\bar{Z}_{s,t},\varnothing_{s,t},p_{s},C)}\|_{2}^{2},$ (6)
$z_{t-1}(\bullet)$ denotes applying DDIM sampling using latent code $\bar{Z}_{s,t}$, source embedding $\varnothing_{s,t}$, source pose $p_{s}$, and text prompt $C$. Building upon the optimized source embeddings $\{\varnothing_{s,t}\}_{t=1}^{T}$ obtained from this process, we then proceed with the second optimization, i.e. pose-aware embedding optimization, where we inject the target pose signals $P=\{p_{i}\}_{i=1}^{N}$ into the optimized pose-aware embeddings {{~$\{\{\widetilde{\varnothing}_{x_{i},t}\}_{t=1}^{T}\}_{i=1}^{N}$, as detailed in Alg. 1. Perceiving the target pose signals, these optimized pose-aware embeddings {{~$\{\{\widetilde{\varnothing}_{x_{i},t}\}_{t=1}^{T}\}_{i=1}^{N}$ ensure a flawless alignment between the generated character actions and the target poses, while upholding the consistency of character-independent content.
@@
Specifically, to incorporate the pose signals, we integrate ControlNet into all processes of the module. Diverging from null-text inversion Mokady et al. (2023) that achieves image reconstruction by optimizing unconditional embeddings Ho and Salimans (2022), our pose-aware inversion optimizes the conditional embedding $\varnothing_{text}$ of the text prompt $C$ during the reconstruction process. The motivation stems from the observation that conditional embedding contains more abundant and robust semantic information, which endows it with a heightened potential for encoding pose signals.
Figure 3: Illustration of Pose-Aware Control Module. Through two optimizations, the pose-aware embeddings are injected with motion awareness, which enables the alignment of generated actions with the target poses while maintaining consistency in character-independent scenes.
@@
1:  Set guidance scale = 1.0. Calculate DDIM inversion Dhariwal and Nichol (2021) latent code $Z_{0},...,Z_{T}$ corresponding to input image $I_{s}$.
2:  Set guidance scale = 7.5. Obtain optimized source embeddings $\{\varnothing_{s,t}\}_{t=1}^{T}$ through pose-aware inversion.